OpenStudy (anonymous):

log2 x = log2 (1-5x)-1 find x

OpenStudy (anonymous):

\[\log_{2} x=\log_{2} (1-5x) - 1 \]

OpenStudy (apoorvk):

1 can be written as log of 2 to the base 2. use properties to simplify and take anti-log.

OpenStudy (anonymous):

@callisto... can help?

OpenStudy (campbell_st):

\[\log _{2} x - \log _{2} ( 1 - 5x) = -1\] raise each term to a power of 2 \[x - (1 - 5x) = 2^{-1}\] 6x - 1 = 1/2 solve the equation

OpenStudy (apoorvk):